You deposit $3,000 in a savings account that earns 6% simple interest. How long will it be before you have a total of $3,500?
Answer by marroqus(3)

Initial saving s=3000
Simple interest i=6% (per year)
Final total F=3500
Time t=?
Equation F-s=s*i*t
t=(F-s)/(s*i)
t=(3500-3000)/(3000*0.06)
t=500/180
t=2.77 years